Borrow or pay cash for an asset Personal Finance Problem   Mark and Stacy McCoy are set to move into their first apartment. They visited Levin Furniture looking for a dining room table and buffet. Dining room sets are typically one of the more expensive home furnishing​ items, and the store offers financing arrangements to customers. Mark and Stacy have the cash to pay for the​ furniture, but it would definitely deplete their savings so they want to look at all their options. The dining room set costs $15,000 and Furniture​ R'Us offers a financing plan that would allow them to either​ (1) put 20% down and finance the balance at 3.8% annual interest over 48 months or​ (2) receive an immediate $500 cash​ rebate, thereby paying only $14,500 cash to buy the furniture. Bob and Carol currently earn 2.8% annual interest on their savings.

a.  Calculate the cash down payment for the loan.

b.  Calculate the monthly payment on the financed amount.  ​(​Hint:Treat the current loan as an annuity and solve for the monthly​ payment.)

c.  Calculate the initial cash outlay under the cash purchase option. Do not forget to reduce the​ cash-price by the cash down payment forgone on the loan​ (part

a​). For parts d​, ​e, and ​f, ignore the time value of money.

d.  Assume that they can earn a simple interest rate of 2.8% on​ savings, what will Bob and Carol give up​ (opportunity cost) over the 4 years if they pay​ cash?

e.  What is the cost of the cash alternative at the end of 4​ years?

f.  Should Bob and Carol choose the financing or the cash​ alternative?

Thank you in advance!
Answer #1

a) Down Payment = $15000 * 20%. = $5,000 b). monthly payment: - Balance Amount = 15000 - 3000 = 12,000 Annual Rate of interes